# Information-Theoretic Measures on Lattices for High-Order Interactions

### v1

@@ Abstract
Traditional models reliant solely on pairwise associations often prove insufficient in capturing the complex statistical structure inherent in multivariate data. Yet existing methods for identifying information shared among groups of $d>3$ variables are often intractable; asymmetric around a target variable; or unable to consider all factorisations of the joint probability distribution. Here, we present a framework that systematically derives high-order measures using lattice and operator function pairs, whereby the lattice captures the algebraic relational structure of the variables and the operator function computes measures over the lattice. We show that many existing information-theoretic high-order measures can be derived by using divergences as operator functions on sublattices of the partition lattice, thus preventing the accurate quantification of all interactions for $d>3$. Similarly, we show that using the KL divergence as the operator function also leads to unwanted cancellation of interactions for $d>3$. To characterise all interactions among $d$ variables, we introduce the Streitberg information defined on the full partition lattice using generalisations of the KL divergence as operator functions. We validate our results numerically on synthetic data, and illustrate the use of the Streitberg information through applications to stock market returns and neural electrophysiology data.
@@ 1 Introduction
Although networks are commonly used as standard descriptions for complex real-world systems, they are inherently restricted to capturing and representing pairwise interactions between units or nodes [1]. Many real-world systems are better characterised by high-order interactions involving groups of three or more units [2, 3, 4, 5]. Yet direct measurements of such group interactions are rarely available, necessitating approaches to reconstruct the systems’ underlying interactions and mechanisms from observed multivariate data [5, 6, 7, 8].
@@
Information theory provides a powerful quantitative framework to dissect complex interactions by detecting high-order interactions with a solid probabilistic basis and the capacity to relate statistical structure to function [9]. Early metrics such as interaction information [10] or co-information [11] began to address these high-order interactions. However, those formulations only consider a subset of the factorisations of the joint probability distribution, thus limiting their scope [12, 13]. More recently, the connected information attempted to capture interactions among $d$ variables by considering their maximum entropy states [14]. Yet, its practical application to real-world data is hampered by the computational difficulties in constructing the maximum entropy distributions [15]. One of the more prevalent measures, Partial Information Decomposition (PID), offers a breakdown of high-order interaction into information atoms [16]. While PID provides a detailed insight into how information is shared among variables, it becomes computationally infeasible for $d>3$ and exhibits asymmetry among variables. The increasing interest in detecting high-order interactions in data and the limitations of existing tools underscore the need for more practical and comprehensive information-theoretic measures that can handle systematically and effectively all factorisations of the joint probability distribution.
@@
Here, we develop a framework that systematically formalises high-order information-theoretic measures using lattice theory. By leveraging lattice embeddings and isomorphisms, we demonstrate that various existing information-theoretic measures can be derived from sublattices of the partition lattice, with a divergence function applied to each element within these lattices. Using the complete partition lattice, we propose a novel information measure, the Streitberg Information, as defined in Equation (6). This measure accounts for all factorisations of the joint probability distribution of $d$ variables, vanishing when a lower-order factorisation is present. We also demonstrate that using Kullback-Leibler (KL) as a divergence function collapses lattice substructures, resulting in undesired, non-trivial equivalence of measures derived from different sublattices. To preserve the desired vanishing property of the partition lattice, we formulate the Streitberg information using a generalisation of the KL divergence, the Tsallis-Alpha divergence, and develop its non-parametric consistent estimator using a $k$-nearest-neighbours method. Finally, we present numerical validations of the Streitberg information on synthetic data, followed by applications to real-world financial and neural electrophysiology data.
2 Information-theoretic measures of interaction
@@
In information theory, given two random variables, $X_{1}$ and $X_{2}$, the amount of information provided about one of the variables given the value of the other can be quantified by the mutual information
$\mathrm{I}(X_{1};X_{2})=\mathrm{H}(X_{1})+\mathrm{H}(X_{2})-\mathrm{H}(X_{1},X_{2})$
@@
where $\mathrm{H}(\cdot)$ is the joint entropy of a set of random variables. The mutual information vanishes if and only if the two variables are independent ($X_{1}\perp\!\!\!\!\perp X_{2}$[17].
@@
The generalisation of mutual information to encompass the dependencies among a set $\mathbf{\mathcal{X}}=\{X_{1},\ldots,X_{d}\}$ of $d>2$ random variables is not trivial. The simplest way to quantify the amount of information shared among $d$ variables is total correlation (TC) [18]:
@@
which vanishes if and only if the variables are jointly independent. Yet $\mathrm{TC}(d)$ fails to provide any details about the exact form of the interactions, i.e., whether it can be explained by the sum of low-order interactions [19, 14]. An alternative extension of mutual information is the interaction information (II) [10] which captures the high-order information in a system of $d$ variables through the inclusion-exclusion principle:
$\mathrm{II}(d)=-\sum_{T\subseteq\mathcal{X}}(-1)^{d-|T|}\,\mathrm{H}(T)$ (2)
where the sum extends over the power set of all subsets $T$ of $\mathcal{X}$ and ||$\cdot$|| ⋅ | denotes the cardinality of the set. Although $\mathrm{II}(d)$ has been widely used, its interpretation when equal to zero is not clear [12]. Below we show that this is because some of the non-trivial high-order terms are decomposed into a sum of lower order terms via KL-divergence (Section 4). We also show that, although seemingly distinct, these generalisations of mutual information to high-orders actually belong to the same family of information-theoretic measures when formalised through lattice theory (Section 3).
@@ 3 Lattice theory
To see how different information-theoretic measures (e.g.,those in Section 2) can be systematically generated from different lattices, we provide a few relevant concepts and results from lattice theory.
A partially ordered set (poset) defined on the set $\mathcal{X}$ with an order $\leq$ is a lattice (denoted by $L$) if the meet (least upper bound) $x\wedge y$ and the join (greatest lower bound) $x\vee y$ exist and are contained in $L$ for all $x,y\in L$. The maximum element and the minimum element of $L$ are denoted as $\hat{\mathbf{1}}$ and $\hat{\mathbf{0}}$, respectively. A subset of $\mathcal{X}$ closed under $\vee$ and $\wedge$ of $L$ forms a sublattice (denoted $\leq$) of $L$. Let $f:\mathcal{X}\rightarrow\mathbb{R}$ be a real-valued function and let $g(\cdot)$ be the sum function of $f(\cdot)$ over the interval $[\hat{\mathbf{0}},y]$. Then we have the following relationships, known as the Möbius inversion theorem [20]:
$g(y)=\sum\zeta(x,y)f(x)\iff f(y)=\sum_{x\leq y}\mu(x,y)g(x),\qquad x,y\in P$ (3)
where the partial order is encoded by the Zeta matrix, with $\zeta(x,y)=1$ if $x\leq y$ and 0 otherwise. Its inverse is the Möbius matrix with elements $\mu(x,y)$. Clearly, a measure derived by Möbius inversion depends on both the structure of the lattice and the choice of $f$ (or $g$); Hence the Möbius inversion on the same poset with different operator functions can lead to different measures. Conversely, in Proposition 1 we show a non-trivial case where different lattices with the same operator function lead to the same (information-theoretic) measure through Möbius inversion.
These links can be understood through lattice homomorphisms [21]. A homomorphism $\varphi$ is a map of the lattice $L_{0}$ into the lattice $L_{1}$ which satisfies
$\varphi(a\vee b)=\varphi(a)\vee\varphi(b),\qquad\varphi(a\wedge b)=\varphi(a)\wedge\varphi(b),\qquad a,b\in L_{0},\quad\varphi(a),\varphi(b)\in L_{1}$
A one-to-one homomorphism is also called an embedding. If $\varphi$ is a bijection, then $\varphi$ is an isomorphism (denoted $L_{0}\cong L_{1}$). Below we show how lattice homomorphisms lead to insights into the relationships between high-order information-theoretic measures.
Figure 1: Lattice isomorphisms and embeddings. a Isomorphisms between all the two-element chains as $d$ increases. b Isomorphism between deatomised $B(3)$ and $L(3)$. Elements in $L(3)$ can be mapped to the non-shaded elements in $B(3)$. $L(3)$ is the same as $P(3)$. c Isomorphism between deatomised $B(4)$ and $L(4)$. $L(4)$ is a sublattice of $P(4)$.
@@
A lattice with $m$ elements is a chain, denoted $C_{m}$, if $a\leq b$ or $b\leq a$ for all $a,b\in C_{m}$. The length of $C_{m}$ is $(m-1)$. See Fig. 1a.
In particular, the two-element chain $C_{2}$ consists only of elements $\hat{\mathbf{1}}$ and $\hat{\mathbf{0}}$ with Möbius inversion
@@
If we make $g$ equal to the Kullback–Leibler (KL) divergence $D_{\operatorname{KL}}(\mathbb{P}\|\mathbb{Q})=\int\mathbb{P}\log(\mathbb{P}/{\mathbb{Q}})\mathrm{d}x$ with respect to $\hat{\mathbf{0}}$, where $\mathbb{P}$ is the probability function, we trivially retrieve the total correlation (1):
$f(\hat{\mathbf{1}})=D_{\operatorname{KL}}\left(\mathbb{P}_{1\cdots d}\|\prod_{i=1}^{d}\mathbb{P}_{i}\right)-D_{\operatorname{KL}}\left(\prod_{i=1}^{d}\mathbb{P}_{i}\|\prod_{i=1}^{d}\mathbb{P}_{i}\right)=\mathrm{TC}(d)$ (5)
Next, the Boolean lattice $B(d)$ (with inclusion partial order) [22] is a natural choice to devise high-order information-theoretic measures as classical information is measured in binary unit bits.
@@
A lattice is distributive if $(x\wedge y)\vee(x\wedge z)=x\wedge(y\vee(x\wedge z))$. A lattice is complemented if there exists $y$ such that $x\wedge y=\hat{\mathbf{1}}$ and $x\vee y=\hat{\mathbf{0}}$. A complemented distributive lattice is a Boolean lattice.
The Boolean lattice is isomorphic to the subset lattice (the powerset of a set ordered by inclusion) and can be used to derive the inclusion-exclusion principle using the Möbius inversion
@@
It then easily follows that using $D_{\operatorname{KL}}$ as operator function $g$ (as defined for $\mathrm{TC}(d)$ in Eq. (5)) on the Boolean lattice $B(d)$ leads to $f(\hat{\mathbf{1}})=\mathrm{II}(d)$ in Eq. (2).
The above lattices are particular embeddings of the partition lattice.
@@
Let $\mathcal{P}$ denote the set of all partitions of $\mathcal{X}$ where a partition $\pi$ is a collection of non-empty, pairwise disjoint subsets (blocks) $b_{j}\subseteq\mathcal{X}$ that cover $\mathcal{X}$. The partition lattice $P(d)$ is defined on the set $\mathcal{P}$ with the refinement ordering, i.e. for $\pi,\sigma\in\mathcal{P}$, $\pi\leq\sigma$ means every block of $\pi$ is contained in a block of $\sigma$.
@@
The two-element chain $C_{2}$ is a sublattice of the partition lattice, $P(d)$, consisting of its minimal and maximal elements, and the chain sublattices for all $d$ are isomorphic (Figure 1a). Hence the $C_{2}$-induced measure $\mathrm{TC}(d)$ only measures the difference between the joint behaviour of $d$ variables and the sum of their individual marginal effects.
Another sublattice of the partition lattice, where partitions contain at most one non-singleton block, is denoted here as the Lancaster lattice $L(d)$ (Figure 1b,c).
@@
For a proof, see Supplementary Section A.2. As a consequence, $B(d)$-based measures can also be generated from $L(d)$, a sublattice of $P(d)$, e.g., the computation of the Lancaster interaction [23, 25].
Remark (Simplicial complexes): A popular representation of high-order systems is through simplicial complexes [29]. The elements in a ($d-1$)-simplex with inclusion ordering form a lattice isomorphic to $B(d)$. Furthermore, the boundary matrices of a ($d-1$)-simplex appear as block submatrices in the Zeta matrix and the Möbius matrix of $P(d)$ (see Supplementary Section C).

### v3

# Information-Theoretic Measures on Lattices for Higher-Order Interactions

@@ Abstract
Traditional measures based solely on pairwise associations often fail to capture the complex statistical structure of multivariate data. Existing approaches for identifying information shared among $d>3$ variables are frequently computationally intractable, asymmetric with respect to a target variable, or unable to account for all the ways in which the joint probability distribution can be factorised. Here we present a systematic framework based on lattice theory to derive higher-order information-theoretic measures for multivariate data. Our construction uses lattice and operator function pairs, whereby an operator function is applied over a lattice that represents the algebraic relationships among variables. We show that many commonly used measures can be derived within this framework, yet they fail to capture all interactions for $d>3$, either because they are defined on restricted sublattices, or because the use of the KL divergence as an operator function, a typical choice, leads to undesired disregard of groups of interactions. To fully characterise all interactions among $d$ variables, we introduce the Streitberg Information, which is defined over the full partition lattice and uses generalised divergences (beyond KL) as operator functions. We validate the Streitberg Information on synthetic data, and illustrate its application in detecting complex interactions among stocks, decoding neural signals, and performing feature selection in machine learning.
@@ 1 Introduction
Understanding the emergent structure of complex systems is a key objective across scientific disciplines, and traditional models built on pairwise interactions have been instrumental in revealing fundamental system effects [1]. However, recent evidence suggests that many real-world systems cannot be adequately described by pairwise interactions alone [2, 3], much like a symphony requires the harmonisation of an entire orchestra, not just two musicians. In biology [4], ecosystems [5], and social networks [6], higher-order interactions (i.e., those involving interactions between groups of more than two elements) are increasingly recognised as critical for driving collective dynamics [2, 3, 5, 7]. Yet direct measurements of such group interactions are rarely available, necessitating approaches to reconstruct the underlying interactions and mechanisms from observed multivariate data [3, 8, 9, 10].
@@
Information theory provides a rigorous mathematical framework to detect statistical interactions between variables with the added capacity to relate statistical structure to function [11]. Early information-based metrics, such as total correlation (TC) [12] and interaction information (II) [13], began to address higher-order interactions inspired by the notion of statistical independence, but their formulations considered only particular factorisations of the joint probability distribution (see Section 3), hence with restricted ability to capture complex interactions between multiple variables [14, 15, 16]. These measures were later followed by connected information (CI), which captured interactions among $d$ variables by considering their maximum entropy states [17]. Yet its practical application to real-world data is hampered by the computational difficulty of constructing the maximum entropy distributions [18]. One of the prevalent measures is partial information decomposition (PID), which offers a breakdown of higher-order interaction into information atoms [19]. While PID details how information is shared among variables, it becomes computationally infeasible for $d>3$ and exhibits asymmetry among variables. The increasing interest in detecting higher-order interactions in data and the limitations of existing tools thus underscore the need for practical information-theoretic measures that can handle systematically such interactions for multivariate data by considering all factorisations of the joint probability distribution.
@@
Our contributions in this work are twofold: (i) First, we develop a mathematical framework that systematically formalises higher-order information-theoretic measures using lattice theory. We demonstrate that many pre-existing information-theoretic measures can be derived from sublattices of the partition lattice with divergences as operator functions over these sublattices. (ii) Second, we propose the Streitberg Information (Equation (6)), a novel information measure based on the complete partition lattice, which accounts for all factorisations of the joint probability distribution of $d$ variables and vanishes when any lower-order factorisation is present. We further demonstrate that, if the Kullback-Leibler (KL) divergence is used as the operator function (a typical choice), it leads to the collapse of lattice substructures, resulting in the undesired, non-trivial cancellation of groups of interactions. Therefore, to avoid this loss of information and to preserve the structure of the full partition lattice, the Streitberg Information (SI) uses the Tsallis-Alpha divergence and we develop its consistent estimator using a $k$-nearest-neighbours method. Finally, we present numerical validation of SI on synthetic data, followed by illustrative applications to finance, neuroscience, and machine learning.
2 Information-Theoretic Measures
@@
In information theory, the shared information between two random variables, $X_{1}$ and $X_{2}$, can be quantified using the mutual information (MI):
$\mathrm{MI}(X_{1};X_{2})=\mathrm{H}(X_{1})+\mathrm{H}(X_{2})-\mathrm{H}(X_{1},X_{2}),$
@@
where $\mathrm{H}(\cdot)$ is the entropy of the random variables. Mutual information vanishes if and only if the two variables are independent [20].
@@
Generalising mutual information to capture dependencies among a set $\mathbf{\mathcal{X}}=\{X_{1},\ldots,X_{d}\}$ of $d>2$ random variables is nontrivial. A straightforward way to quantify the total shared information among $d$ variables is through total correlation (TC) [12]:
@@
which vanishes if and only if the variables are jointly independent. However, $\mathrm{TC}(d)$ provides limited insight into the specific nature of the interactions that may exist between variables  [17, 21]. An alternative extension of MI is the interaction information (II) [13], which captures the higher-order information in a system of $d$ variables through the inclusion-exclusion principle:
$\mathrm{II}(d)=-\sum_{T\subseteq\mathcal{X}}(-1)^{d-|T|}\,\mathrm{H}(T),$ (2)
where the sum extends over the power set of all subsets $T$ of $\mathcal{X}$ and $|T|$ denotes the cardinality of set $T$. Although widely used, the interpretation of the vanishing of $\mathrm{II}(d)$ is not clear [14]. Indeed, we show below that this is because non-trivial higher-order terms are decomposed into a sum of lower-order terms via KL divergence (Section 4). We also show below that, although seemingly distinct, both TC and II actually belong to the same family of information-theoretic measures when formalised through lattice theory (Section 3).
@@ 3 Lattice Theory
To understand how different information-theoretic measures can be systematically generated from different lattices, we first provide a few relevant concepts and results from lattice theory.
A partially ordered set (poset) defined on $\mathcal{X}$ with partial order $\leq$ is a lattice (denoted by $\mathcal{L}$) if any two elements $x,y\in\mathcal{L}$ have a unique supremum $x\vee y$ (called join) and a unique infimum $x\wedge y$ (called meet). The maximum element and the minimum element of $\mathcal{L}$ are denoted as $\hat{\mathbf{1}}$ and $\hat{\mathbf{0}}$, respectively. A subset of $\mathcal{L}$ that is closed under $\vee$ and $\wedge$ forms a sublattice $\mathcal{S}$ of $\mathcal{L}$ (denoted $\mathcal{S}\leq\mathcal{L}$).
Let $f:\mathcal{X}\rightarrow\mathbb{R}$ be an operator function and let $g(\cdot)$ be the sum function of $f(\cdot)$ over the interval $[\hat{\mathbf{0}},y]$. Then we have the following relationship, known as the Möbius inversion theorem [22, 23]:
$g(y)=\sum\zeta(x,y)f(x)\iff f(y)=\sum_{x\leq y}\mu(x,y)g(x),$ (3)
where the partial order is encoded by the Zeta matrix with elements $\zeta(x,y)=1$ if $x\leq y$ and 0 otherwise, $\forall x,y\in\mathcal{L}$. The inverse of $\zeta(x,y)$ is the Möbius matrix with elements $\mu(x,y)$.
Clearly, a measure derived by Möbius inversion depends on both the structure of the lattice and the choice of $f$ (or $g$). Hence, Möbius inversion on the same lattice with different operator functions can lead to different measures. Conversely, in Proposition 1 we show a non-trivial case where Möbius inversion on different lattices with the same operator function leads to the same (information-theoretic) measure.
Figure 1: Lattice embeddings. The black dots indicate the marginal distributions of the singletons ($p_{i}$). The line, triangular and square shapes represent the joint distribution of two, three and four variables, respectively. (a) The two-element chain is isomorphic for all $d$. (b) The deatomised $B(3)$ is isomorphic to $L(3)$, and elements in $L(3)$ can be mapped to the non-shaded elements in $B(3)$. $L(3)$ is a sublattice of $P(3)$ (in this case equal). (c) For $d=4$, we see that the deatomised $B(4)$ is isomorphic to $L(4)$, which is a (strictly smaller) sublattice of $P(4)$.
To clarify the links between different lattices it is useful to consider lattice homomorphisms [24]. A homomorphism $\varphi$ is a map of the lattice $\mathcal{L}_{0}$ into the lattice $\mathcal{L}_{1}$ which satisfies
$\varphi(a\vee b)=\varphi(a)\vee\varphi(b),\qquad\varphi(a\wedge b)=\varphi(a)\wedge\varphi(b)$
where $a,b\in\mathcal{L}_{0}$ and $\varphi(a),\varphi(b)\in\mathcal{L}_{1}$. A one-to-one homomorphism is also called an embedding. If $\varphi$ is a bijection, then $\varphi$ is an isomorphism (denoted $\mathcal{L}_{0}\cong\mathcal{L}_{1}$).
Now we introduce briefly the main types of lattices that underpin the definitions of different higher-order information-theoretic measures and highlight their relationships via lattice homomorphisms.
First, we consider the simplest lattice, the chain, which underpins the notion of total correlation.
@@
A lattice with $m$ elements is a chain, denoted $C_{m}$, if $a\leq b$ or $b\leq a$ for all $a,b\in C_{m}$, i.e. all the elements in $C_{m}$ are directly comparable. The length of $C_{m}$ is $(m-1)$.
The simplest such lattice is the two-element chain $C_{2}$, consisting of elements $\hat{\mathbf{1}}$ and $\hat{\mathbf{0}}$ with Möbius inversion
@@
Let $\hat{\mathbf{1}}$ and $\hat{\mathbf{0}}$ be the joint probability density function $p_{X_{1},\ldots,X_{d}}=:p_{1\cdots d}$ and the complete factorisation $\Pi^{d}_{i=1}p_{X_{i}}=:\Pi^{d}_{i=1}p_{i}$, respectively, and let us choose $g(\cdot)=D_{\operatorname{KL}}\left(\cdot\,\|\prod_{i=1}^{d}p_{i}\right)$, where $D_{\operatorname{KL}}(p\|q)=\int p\,\log(p/{q})\mathrm{d}x$ is the Kullback–Leibler (KL) divergence between the probability density functions $p$ and $q$. Then Equation (4) becomes:
$f(\hat{\mathbf{1}})=$ $D_{\operatorname{KL}}\left(p_{1\cdots d}\|\prod_{i=1}^{d}p_{i}\right)-D_{\operatorname{KL}}\left(\prod_{i=1}^{d}p_{i}\|\prod_{i=1}^{d}p_{i}\right)$
$=$ $D_{\operatorname{KL}}\left(p_{1\cdots d}\|\prod_{i=1}^{d}p_{i}\right)-0=\mathrm{TC}(d),$ (5)
thus trivially recovering the total correlation $\mathrm{TC}(d)$ in Equation (1).
Next, we consider the Boolean lattice $B(d)$ defined with the inclusion partial order [25]. $B(d)$ is a natural choice for higher-order information-theoretic measures as classical information is measured in binary unit bits.
@@
For a positive integer $d$, the Boolean lattice of order $d$, denoted by $B(d)$, is the poset on all the subsets of $[d]={1,\ldots,d}$ equipped with the inclusion relation.
The Boolean lattice is isomorphic to the subset lattice (the power set of a set ordered by inclusion) and the inclusion-exclusion principle appears naturally in this lattice through the Möbius inversion
@@
If we choose $g(\cdot)=D_{\operatorname{KL}}\left(\cdot\,\|\prod_{i=1}^{d}p_{i}\right)$ as an operator function (as shown in Equation (5)) on the Boolean lattice $B(d)$, then it follows easily that $f(\hat{\mathbf{1}})=\mathrm{II}(d)$, i.e., we recover the interaction information in Equation (2).
The above lattices are embeddings of the partition lattice, which plays a central role in our derivations.
@@
Let $\mathcal{P}$ denote the set of all partitions of $\mathcal{X}$, where a partition $\pi$ is a collection of non-empty, pairwise disjoint subsets (blocks) $b_{j}\subseteq\mathcal{X}$ that cover $\mathcal{X}$. The partition lattice $P(d)$ is defined on $\mathcal{P}$ with the refinement ordering, i.e. for $\pi,\sigma\in\mathcal{P}$, $\pi\leq\sigma$ means every block of $\pi$ is contained in a block of $\sigma$.
@@
The minimal ($1|2|\cdots|d$) and maximal ($12\cdots d$) elements in $P(d)$ form a sublattice denoted as $C_{2}(d)$, which is isomorphic to the two-element chain $C_{2}$ for any $d$ (Figure 1a). Hence the $C_{2}$-induced measure $\mathrm{TC}(d)$ always measures the difference between the top of the lattice, $\hat{1}$, and the bottom of the lattice, $\hat{0}$, i.e., between the joint distribution and the product of all the marginals, independently of $d$. As $d$ grows, the lattice $C_{2}(d)$ remains static (just two elements) and $\mathrm{TC}$ becomes progressively less informative.
Another sublattice of the partition lattice that is of interest here is the sublattice where partitions contain at most one non-singleton block. We denote this sublattice as the Lancaster lattice $L(d)$ (see Figure 1b,c).
@@
Proof.
For a proof, see Appendix A.1. ∎
As a consequence, $B(d)$-based measures can also be generated from $L(d)$, a sublattice of $P(d)$, e.g., the computation of the Lancaster interaction [26, 16].
Remark (Simplicial complexes): A popular representation of higher-order systems is through simplicial complexes [31]. The elements in a ($d-1$)-simplex with inclusion ordering form a lattice isomorphic to the Boolean lattice $B(d)$. Furthermore, the boundary matrices of a ($d-1$)-simplex appear as block submatrices in the Zeta matrix and the Möbius matrix of the partition lattice $P(d)$ (see Appendix B).